OpenStudy (anonymous):

I am really frustrated-no one seems to want to help--Again the problem is v(x)=x^3-5x^2-2x+24. The depth is x+2. I am needing to know the linear expressions with integer co-efficients for the other dimensions. AND if the depth of the chest is 75 feet, what are the other dimensions? Please provide the answers and show all work as I am TOTALLY CONFUSED

OpenStudy (turingtest):

V(x) is some function that represents the volume I presume?

OpenStudy (turingtest):

the volume of you box, like any, is\[V=\ell wh\]only in this case the catch is that the length, width, and height are all given by polynomials

OpenStudy (amistre64):

http://openstudy.com/users/kf#/updates/4fb3d8e4e4b055653429a741 noone wants to help you out?

OpenStudy (turingtest):

we know that\[h=x+2\]so we can then figure that\[\frac Vh={x^3-5x^2-2x+24\over x+2}=lw\]doing the polynomial ling division will get you a quadratic expression, which I'm betting can be factored

OpenStudy (anonymous):

amistre64 I had two people but they had to go-I am really stumped here.

OpenStudy (turingtest):

ok I am one do you understand what I just said?

OpenStudy (anonymous):

Turing Test-I do understand what you are saying-you were not one of the two who were helping on this particular problem earlier

OpenStudy (turingtest):

I know that here is a different explanation do you prefer that I continue on your other thread from where the others left off?

OpenStudy (turingtest):

if you do understand what I;m saying, use polynomial long division to get an expression for \[\frac Vh\]

OpenStudy (anonymous):

thanks so much

OpenStudy (turingtest):

..or synthetic division if you know it

OpenStudy (turingtest):

once you have simplified \[\frac Vh\] you will be left with a quadratic that can be factored

OpenStudy (anonymous):

Thanks for helping out I really do appreciate it

OpenStudy (turingtest):

you are welcome, but I'd like to know that you are getting the right answer... did you try the division?